# Botender: Supporting Communities in Collaboratively Designing AI Agents through Case-Based Provocations

### v1

@@ 5.3. Validation Study Results
Figure 8. Participants’ ratings on various aspects of cases generated by Botender’s case-based provocation algorithm and the baseline algorithm. The figure shows notched box plots, with the notches indicating the medians, and the means and 95% confidence intervals overlaid on the plots. The results show significant differences for all rated aspects.:
@@ 6. Field Study
To understand how people use Botender as an end-to-end system for collaborative AI agent design, we conducted a field study in real-world Discord communities. On Discord, each server is an online community where individuals with shared interests connect and interact. These interests range from casual hobbies like gaming and anime to professional topics such as programming or specialized spaces for customer service (Aquino et al., 2025). Our goal in this field study was to understand how diverse Discord communities use Botender to collaboratively design and customize bots for their servers.
@@ 6.1. Study Procedures
The field study lasted five days for each group of participants, with each group consisting of about five people from the same Discord servers who knew each other well. All participants were active community members or admins in their servers, making them ideal candidates to design their own community bots using Botender.
The study began with a synchronous onboarding session, where participants received guidance on how to use Botender and had the chance to try the system in a dedicated onboarding server. We also recorded a comprehensive system walkthrough for a few participants who were unable to attend the group onboarding session. After onboarding, participants installed Botender on their own Discord server and selected the start date for the five-day study period. We set two minimum participation requirements for the study. First, each participant needed to create or edit at least one proposal per day. Secondly, as a group, they were required to deploy at least three tasks tailored to their community’s specific needs, norms, and values by the end of the study. We kept the participation requirements minimal to provide them with the flexibility to decide when and how much they want to engage in bot design, while also ensuring that participants would have ample opportunities for interaction across the field study period (cf. (Zhang and Cranshaw, 2018; Kuo et al., 2025)). We encouraged participants to design tasks that reflected their community’s unique norms and culture, rather than purely logistical tasks.
@@
6.2. Recruitment
Group ID Group Size Server Size Server Type
G1 5 5 A close-knit friend group server for hanging out and having fun
We recruited six participant groups via social media, paper flyers, and word of mouth. These groups represented a diverse array of Discord communities, ranging from close-knit friend groups and student organizations to the fan community of an indie band. Each participant received \$100 USD for the field study [^13] During the study, other members of each server who were not participants did not use Botender to design bots., consistent with compensation provided in previous week-long HCI research (cf. (Zhang and Cranshaw, 2018; Kuo et al., 2024)). Table 1 provides the details about each participant group. [^14] In the field study results, participant IDs are indicated with an S, and the groups they belong to are indicated with a G.
@@ 6.3. Study Results
We present the findings from our field study in the following sections. Section 6.3.1 presents the types of tasks participants designed for their community bots and their perceptions of these tasks, as shared in the post-study survey. Our results show that participants were able to design a wide variety of tasks tailored to their communities, and they felt these tasks reflected their unique community needs and culture. Section 6.3.2 explores how participants collaborated using case-based provocations. Across the six communities, participants created a total of over 100 proposals and 800 saved test cases during the study. Analysis of the multiple-choice questions that participants answered after each proposal edit revealed they were more likely to iterate on proposals in response to Botender’s case-based provocations, which encouraged collective reflection and discussion about desirable bot behaviors. Feedback from the post-study survey provides further insight into how participants collaboratively improved and discussed bot designs based on opportunities identified through these cases. Finally, Section 6.3.3 presents additional feedback from participants on aspects of Botender’s design, beyond the case-based provocations, that they found particularly helpful for collaborative bot design. These include the overall system workflow, seamless integration with the community platform, and the use of natural language for bot design. Overall, participants found that the bot behaved in ways that aligned with their needs and community norms, attributing this to the collaborative design process enabled by Botender. Over 97% of participants expressed interest in continuing to use Botender after the study period.
@@
Table 2 presents descriptive statistics from the field study, including the number of tasks each group deployed and the number of proposals and cases they created to support task deployment. To provide a glimpse of the resulting tasks, Table 3 presents a sample of tasks deployed by each group. As one participant noted, the range of tasks is quite broad and diverse, spanning everything “from funny quips to actual advice and help to games and welcome messages” (S4, G5). The tasks participants created for the bot align with prior research (Seering et al., 2018), demonstrating a balance between task-oriented and socially-oriented functions, or combinations of both. As shown in Table 3, the types of tasks participants created generally reflect the nature of the server, whether it is more professionally or socially oriented.
@@
Figure 9. The percentage of proposal edits aimed at addressing specific cases participants reviewed versus general issues that they thought of during the proposal iteration process. A significantly higher proportion of edits were inspired by specific cases.:
@@
Finally, participants found the test cases helpful in guiding their collective decision-making on whether to deploy a proposal. For example, as one participant shared: “It’s very easy for a chatbot to misunderstand a prompt, or tack on unnecessary things, the same goes for a human when only looking at what the prompt says, and the actual test cases gives a realistic view of what the implementation will look like, what Botender will/could say, and won’t/couldn’t say. We had a task that referenced religion, which was a controversial topic to some, but the test cases served to show that Botender was doing so in an appropriate way and did not step out of line, and that eventually made us agree on implementing the task” (S15, G4). Similarly, another participant shared that the test cases sometimes led them to decide against deploying a proposal after seeing how the bot would actually respond: “I saw this with the gaslighting bot, when people in the group saw the responses it generated were in general invalidating of ones feelings, the group was hesitant on approving it” (S18, G1). Participants found the test cases useful for collective decision-making because “the test cases give us a good idea on how the bot reacts to different situations, it also shows where it thrives/where its more limited in its capabilities” (S13, G5).
@@
Participants also emphasized that deep integration with the community platform they already use is key to keeping community members actively engaged in the bot design process. They agreed that “the experience of designing the bot with the Botender system integrates well with their usage of Discord” ($M=6.14,SD=1.57$), and shared specific integrations they found particularly valuable. For example, many participants expressed how much they appreciated having a dedicated discussion thread for each proposal on Discord: “It was an amazing user experience, having it automatically create threads and open/close proposals automatically was very efficient, and felt very modern” (S11, G6). They found it “really cool that it gives live updates via threads on discord, that was a great idea that is actually useful” (S13, G5) and believed that “the built-in notifications kept everyone in the loop. For example, whenever someone made a new proposal, the notification feature made it easy to see updates right away and give feedback without missing anything” (S2, G3). They found this approach effectively coordinated discussions because “using threads for the ticketing system eliminates need to manually organize the discussions and keeps everything in one place” (S6, G5), and noted that “the use of threads within discord for the discussion of proposals is the best way it could be done in my opinion” (S25, G4). In addition to the threading design, participants found “the UI and layout of the [Botender] website is extremely intuitive ” (S26, G4) and appreciated how details like the login process and overall aesthetics seamlessly fit with their use of Discord. For example, participants shared “it’s very simple to manage the bot via the OAuth login dashboard and we had no issues because of its deep integration with Discord” (S6, G5), and “I think this is my favorite part about Botender! […] it is very easy to use and very much fits the vibes and general aesthetic of Discord very well” (S15, G4). Overall, participants found “Botender fit naturally into how I already use Discord […], which made collaboration smooth and kept the focus on shaping the bot, not juggling tools” (S2, G3).
@@
Finally, while using natural language to write prompts is not a unique feature of the Botender system, several participants noted that this greatly facilitates broader community participation in bot design. For example, a participant shared that “the way it’s designed allows me to describe tasks in plain English, and it understands what I mean without needing technical setup. […] It reduces friction, lets me focus on shaping the community’s culture, and makes it easier to adapt the bot’s behavior to fit our needs. […] That makes it accessible not only to me but also to other community members who may not have a technical background” (S2, G3). A participant without coding background echoed this sentiment, sharing that setting up other bots was much more difficult compared to Botender: “We liked how easy it was to implement ideas. We could talk about them, and then up vote or down vote them based on what we thought. In the past anything like this was really difficult to use. It took a lot of planning and stress. This did not feel that way at all” (S27, G2). Overall, participants found “the bot is a great addition to the server. The fact it’s so easy to customize makes it really simple for any member to contribute, and shape something that reflects what we needs” (S25, G4). In the post-study survey, 97% of participants expressed interest in continuing to use Botender after the study.
@@ 7. Discussion
Bots play vital roles and act as essential socio-technical infrastructure within online communities. It is crucial for communities to collaboratively design bots that meet their specific needs and norms, rather than leaving this to outsiders or just a few technically skilled members. In this paper, we present Botender, a system that supports communities in collaboratively proposing, iterating on, and deploying bots powered by LLM-based agents. In particular, Botender facilitates this collaborative design process through case-based provocations, concrete interaction scenarios generated to provoke user reflection and discussion about desirable bot behaviors within their community. We conducted a validation study and a field study to understand how people perceive and use these case-based provocations to collaboratively design bots in real-world communities. Through a validation study (Section 5), we saw that participants found Botender’s case-based provocations revealed more opportunities for bot improvement, compared with standard test cases. Through a field study (Section 6), we found that real-world Discord communities effectively collaborated on designing bots tailored to their specific needs and norms using Botender’s case-based provocations. This collaboration was further supported by Botender’s overall collaborative workflow, seamless integration with the community platform, and the use of natural language for bot design. In this section, we discuss future directions for HCI systems to better support the collaborative and participatory design of community bots powered by AI agents.
@@ 7.1. Expanding the Capabilities of AI Agents
As the first attempt to support collaborative design of AI agents in community contexts, the current version of Botender focuses on enabling the design of single-turn, LLM-based conversational AI agents. This means that the bot and its underlying agents can process only one user message at a time and respond with a single message. Channel names are the only context information that allows the agents to tailor their behavior more specifically. However, as mentioned in Section 4.5, Botender’s system and agent architecture is designed to handle a broader range of platform events, actions, and context information. On the event side, in addition to listening to user messages, Botender’s listener can detect a wider range of platform events provided by the platform API, such as users joining channels, the creation of new threads, or changes to user permissions. These events can be translated into information that agents can interpret and use to provide appropriate action instructions. Similarly, on the action side, Botender’s action executor can carry out a broader range of actions available through the API, such as creating new channels, muting users when appropriate, or searching the internet for up-to-date information. [^15] We experimented with internet search, but found that it made case generation too slow and affected the user experience. It may become faster as LLMs improve. The context information available to agents can also be expanded depending on the event. For example, for the event of a user sending a message, the context could include the time, the user’s permission, or ten previous messages, providing the agent with richer context to guide its behavior. Future systems could expand agent capabilities based on what would be most useful to communities. However, as prior research points out (Hwang et al., 2024), granting bots more permissions can inevitably raise concerns among community members about potential misuse and the risk of unrecoverable consequences. Striking the right balance between agent capabilities and user concerns will be an important challenge to address.
7.2. Advancing Case-Based Provocations
We see several opportunities to further explore the concept of case-based provocations for prompt and bot design. First, the case-based provocation algorithm presented in this paper represents a proof-of-concept. We see an exciting space for future research to explore the design of case generation algorithms that can more effectively provoke user reflection on opportunities for prompt improvement and surface points for discussion in collaborative design settings. For example, future work might leverage data from users’ collaborative interactions (e.g., edits, discussions, and votes on previously displayed cases) to guide the generation of cases with potential to surface community disagreements. Furthermore, the case-based provocations in our study focused on single-turn interactions. It remains an open question how to generate effective case-based provocations for more advanced bots that respond based on longer message contexts (i.e., multi-turn interactions among a group of people), while still ensuring that generated cases are concise enough to support rapid review in the midst of a collaborative design process. Finally, as agents’ capabilities expand, it will be necessary to advance the case-based provocation algorithm accordingly. Consider a simple functional expansion in which the agents not only read user messages but also examine images sent to the group for content moderation purposes. In this scenario, case-based provocations would consist of a wide variety of generated images created to provoke user reflection and discussion about which types of images should be moderated. As the system expands further to support a broader range of platform events and actions, the complexity of provocations, and the algorithms required to generate them, will increase significantly. Future research will need to explore what these provocations might look like and how algorithms can most effectively generate them to support user reflection and discussion on desirable agent behavior.

### v2

@@ 5.3. Study Results
Figure 8. Participants’ ratings on various aspects of cases generated by Botender’s case-based provocation algorithm and the baseline algorithm. The figure shows notched box plots, with the notches indicating the medians, and the means and 95% confidence intervals overlaid on the plots. The results show significant differences for all rated aspects.
A point range chart displaying the mean and 95% confidence intervals for participants’ ratings on four aspects of the cases (provocativeness, controversialness, coverage, and diversity) generated by Botender’s case-based provocation algorithm compared to the baseline algorithm.
@@ 6. Field Study
To understand how people use Botender as an end-to-end system for collaborative AI agent design, we conducted a 5-day field study in real-world Discord communities. On Discord, each server is an online community where individuals with shared interests connect and interact. These interests range from casual hobbies like gaming and anime to professional topics such as programming or specialized spaces for customer service (Aquino et al., 2025). Our goal in this field study was to understand how diverse Discord communities use Botender to collaboratively design and customize bots for their servers.
6.1. Recruitment
Group ID Group Size Server Size Server Type
G1 5 5 A close-knit friend group server for hanging out and having fun
We recruited six participant groups via social media, paper flyers, and word of mouth. These groups represented a diverse array of Discord communities, ranging from close-knit friend groups and student organizations to the fan community of an indie band. Table 1 provides details about each participant group. [^14] During the study, other members of each server who were not participants did not use Botender to design bots. All participants were active community members or admins in their servers, making them ideal candidates to design their own community bots using Botender. Each participant received \$100 USD for the field study, consistent with compensation provided in previous week-long HCI research (cf. (Zhang and Cranshaw, 2018; Kuo et al., 2024)).
@@ 6.2. Study Procedures
The field study lasted five days for each group of participants. The study began with a synchronous onboarding session, where participants received guidance on how to use Botender and had the chance to try the system in a dedicated onboarding server. We also recorded a comprehensive system walkthrough for a few participants who were unable to attend the group onboarding session. After onboarding, participants installed Botender on their own Discord server and selected the start date for the five-day study period. We set two minimum participation requirements for the study. First, each participant needed to create or edit at least one proposal per day. Secondly, as a group, they were required to deploy at least three tasks tailored to their community’s specific needs, norms, and values by the end of the study. We kept the participation requirements minimal to provide them with the flexibility to decide when and how much they want to engage in bot design, while also ensuring that participants would have ample opportunities for interaction across the field study period (cf. (Zhang and Cranshaw, 2018; Kuo et al., 2025)). We encouraged participants to design tasks that reflected their community’s unique norms and culture, rather than purely logistical tasks.
@@ 6.3. Study Results
We present the findings from our field study in the following sections. [^15] In the field study results, participant IDs are indicated with an S, and the groups they belong to are indicated with a G. Section 6.3.1 presents the types of tasks participants designed for their community bots and their perceptions of these tasks, as shared in the post-study survey. Our results show that participants were able to design a wide variety of tasks tailored to their communities, and they felt these tasks reflected their unique community needs and culture. Section 6.3.2 explores how participants collaborated using case-based provocations. Across the six communities, participants created a total of over 100 proposals and 800 saved test cases during the study. Analysis of the multiple-choice questions that participants answered after each proposal edit revealed they were more likely to iterate on proposals in response to Botender’s case-based provocations, which encouraged collective reflection and discussion about desirable bot behaviors. Feedback from the post-study survey provides further insight into how participants collaboratively improved and discussed bot designs based on opportunities identified through these cases. Finally, Section 6.3.3 presents additional feedback from participants on aspects of Botender’s design, beyond the case-based provocations, that they found particularly helpful for collaborative bot design. These include the overall system workflow, seamless integration with the community platform, and the use of natural language for bot design. Overall, participants found that the bot behaved in ways that aligned with their needs and community norms, attributing this to the collaborative design process enabled by Botender. Over 96% of participants expressed interest in continuing to use Botender after the study period.
@@
Table 2 presents descriptive statistics from the field study, including the number of tasks each group deployed and the number of proposals and cases they created to support task deployment. To provide a glimpse of the resulting tasks, Table 3 presents a sample of tasks deployed by each group. While the current version of Botender only supports the creation of tasks involving single-turn interactions (as mentioned in Section 4.1), the range of tasks is quite broad and diverse. As one participant noted, these tasks span everything “from funny quips to actual advice and help to games and welcome messages” (S4, G5). The tasks participants created for the bot align with prior research (Seering et al., 2018), demonstrating a balance between task-oriented and socially-oriented functions, or combinations of both. As shown in Table 3, the types of tasks participants created generally reflect the nature of the server, whether it is more professionally or socially oriented.
@@
Figure 9. The percentage of proposal edits aimed at addressing specific cases participants reviewed versus general issues that they thought of during the proposal iteration process. A significantly higher proportion of edits were inspired by specific cases.
A bar chart illustrating the percentage of proposal edits made to address specific cases reviewed by participants versus general issues identified during the proposal iteration process. Edits targeting specific cases accounted for 59%, while those addressing general issues made up 37%.
@@
Finally, participants found the test cases helpful in guiding their collective decision-making on whether to deploy a proposal. For example, as one participant shared: “It’s very easy for a chatbot to misunderstand a prompt, the same goes for a human when only looking at what the prompt says. The actual test cases gives a realistic view of what the implementation will look like. We had a task that referenced religion, which was a controversial topic to some, but the test cases served to show that Botender was doing so in an appropriate way and did not step out of line, and that eventually made us agree on implementing the task” (S15, G4). Similarly, another participant shared that the test cases sometimes led them to decide against deploying a proposal after seeing how the bot would actually respond: “I saw this with the gaslighting bot, when people in the group saw the responses it generated were in general invalidating of ones feelings, the group was hesitant on approving it” (S18, G1). Participants found the test cases useful for collective decision-making because “the test cases give us a good idea on how the bot reacts to different situations, it also shows where it thrives/where its more limited in its capabilities” (S13, G5).
@@
Participants also emphasized that deep integration with the community platform they already use is key to keeping community members actively engaged in the bot design process. They agreed that “the experience of designing the bot with the Botender system integrates well with their usage of Discord” ($M=6.14,SD=1.57$), and shared specific integrations they found particularly valuable. For example, many participants expressed how much they appreciated having a dedicated discussion thread for each proposal on Discord: “It was an amazing user experience, having it automatically create threads and open/close proposals automatically was very efficient, and felt very modern” (S11, G6). They believed that “the built-in notifications kept everyone in the loop. For example, whenever someone made a new proposal, the notification feature made it easy to see updates right away and give feedback without missing anything” (S2, G3). They found this approach effectively coordinated discussions because “using threads for the ticketing system eliminates need to manually organize the discussions and keeps everything in one place” (S6, G5). In addition to the threading design, participants appreciated how details like the login process and overall aesthetics seamlessly fit with their use of Discord. For example, participants shared “it’s very simple to manage the bot via the OAuth login dashboard and we had no issues because of its deep integration with Discord” (S6, G5), and “I think this is my favorite part about Botender! […] it is very easy to use and very much fits the vibes and general aesthetic of Discord very well” (S15, G4). Overall, participants found “Botender fit naturally into how I already use Discord […], which made collaboration smooth and kept the focus on shaping the bot, not juggling tools” (S2, G3).
@@
Finally, while using natural language to write prompts is not a unique feature of the Botender system, several participants noted that this greatly facilitates broader community participation in bot design. For example, a participant shared that “the way it’s designed allows me to describe tasks in plain English, and it understands what I mean without needing technical setup. […] It reduces friction, lets me focus on shaping the community’s culture, and makes it easier to adapt the bot’s behavior to fit our needs. […] That makes it accessible not only to me but also to other community members who may not have a technical background” (S2, G3). A participant without coding background echoed this sentiment, sharing that setting up other bots was much more difficult compared to Botender: “We liked how easy it was to implement ideas. We could talk about them, and then up vote or down vote them based on what we thought. In the past anything like this was really difficult to use. It took a lot of planning and stress. This did not feel that way at all” (S27, G2). Overall, participants found “the bot is a great addition to the server. The fact it’s so easy to customize makes it really simple for any member to contribute, and shape something that reflects what we needs” (S25, G4). In the post-study survey, 96% of participants (30 out of 31) expressed interest in continuing to use Botender after the study.
@@ 7. Discussion
Bots play vital roles and act as essential socio-technical infrastructure within online communities. It is crucial for communities to collaboratively design bots that meet their specific needs and norms, rather than leaving this to outsiders or just a few technically skilled members. In this paper, we present Botender, a system that supports communities in collaboratively proposing, iterating on, and deploying bots powered by LLM-based agents. In particular, Botender facilitates this collaborative design process through case-based provocations, concrete interaction scenarios generated to provoke user reflection and discussion about desirable bot behaviors within their community. We conducted a validation study and a field study to understand how people perceive and use these case-based provocations to collaboratively design bots in real-world communities. Through a validation study (Section 5), we saw that participants found Botender’s case-based provocations revealed more opportunities for bot improvement, compared with standard test cases. Through a field study (Section 6), we found that real-world Discord communities effectively collaborated on designing bots tailored to their specific needs and norms using Botender’s case-based provocations. This collaboration was further supported by Botender’s overall collaborative workflow, seamless integration with the community platform, and the use of natural language for bot design.
In this section, we discuss limitations of the current instantiation of Botender that participants wished to overcome, as expressed in their post-study surveys or in their unsuccessful attempts to design bots during the study. For each limitation, we outline opportunities for future HCI research to better support the collaborative and participatory design of community bots powered by AI agents.
7.1. Advancing Case-Based Provocations
7.1.1. Limitations: The current case-based provocation algorithm does not highlight potential conflicts between tasks, and may require more inputs to better surface potential disagreements.
The case-based provocation algorithm presented in this paper represents a proof-of-concept. Our current case-based provocation algorithm generates three specific types of cases, corresponding to common prompt design pitfalls, such as ambiguous phrasing within prompts, which also represent potential areas of disagreement among collaborators. While most participants found that the current case-based provocations effectively facilitated iterative, collaborative design, they also identified opportunities for improvement. For example, participants expressed a desire to see provocations that reveal situations where two tasks might be in conflict, in order to help “prevent conflicting task triggers” (S18, G1). As another example discussed in Section 6.3.2, one participant group (G2) didn’t encounter much disagreement “because we all had similar thoughts” (S27, G2) and “everybody seemed to be on the same page” (S29, G2). While this may simply reflect the dynamics of that particular group, it could also indicate opportunities to improve the case generation algorithm, as we discuss below.
7.1.2. Opportunities: Future algorithms could generate additional types of case-based provocations and incorporate user feedback on cases to better identify potential areas of disagreement.
There are several opportunities for future research to explore the design of case generation algorithms that can more effectively provoke user reflection and discussion throughout iterative design processes. For example, building on participants’ feedback, future work could explore new types of cases not covered in this study, such as those arising from other common prompt design pitfalls, cases that highlight conflicting tasks, or cases that are at the borderline of acceptability for a specific community. We see potential for future case-based provocation algorithms to more effectively surface potential disagreements within a community by learning about divergent perspectives within the community over time, based on users’ interactions—such as users’ thumbs up/down feedback on individual cases or their discussions about bot behavior (cf. (Koshy et al., 2025)). More broadly, we see exciting opportunities to extend the concept of case-based provocations beyond prompt design. For example, future research could explore how algorithmically generated cases might be used to provoke reflection, deliberation, and improvements in the design of other kinds of interactive technologies, or the development of public policy (Kuo et al., 2025).
7.2. Enhancing Support for Prompt Refinement
7.2.1. Limitations: Some participants desired more direct assistance in refining prompts
With the current version of Botender, users can reflect on and discuss a case generated from a prompt, and then manually refine the prompt as needed. However, some participants anticipated that giving a thumbs up or down to a case would not only surface disagreements among users, but also directly contribute to improving the bot. For example, a participant initially “expected the bot will learn [from] it and change the response to something else, but it didn’t” (S5, G3). Relatedly, participants expressed a desire to manually correct a case by entering “the expected output to make the bot learn” (S5, G3), or to “directly explain to the bot why a case is good and why a case is bad, so it would be able to improve” (S7, G5). This feedback highlights a desire for more assistance in prompt refinement, potentially based on users’ feedback on cases, rather than leaving the refinement process completely manual.
7.2.2. Opportunities: Future systems could offer prompt refinement suggestions based on a variety of human feedback on cases.
Building upon participants’ feedback, future systems could consider offering more ways for users to provide feedback on cases, such as allowing them to correct the bot’s response or provide direct explanations (McDaniel and Myers, 1999), in addition to giving a thumbs up or down. The system could then incorporate this human feedback to suggest refined prompts, which users would have the agency to adopt, revise, or ignore as they see fit. While a fully automated prompt refinement process, where the bot evolves on its own without users having access to its prompt, may seem simpler and perhaps more desirable, prior HCI research suggests otherwise (Smith et al., 2020). Communities often prefer to retain agency and control over the AI systems they use or rely on, rather than having this control completely hidden within the system or managed solely by outsiders (Kuo et al., 2024; Halfaker and Geiger, 2020). Moreover, communities’ needs and preferences may change over time, meaning that objectives set in the past may no longer align with their current goals (Halfaker et al., 2025). For these reasons, we recommend that future work seek a balance, to keep the community actively involved and at the center of the design process.
@@ 7.3. Expanding the Capabilities of AI Agents
7.3.1. Limitations: Participants wanted to customize tasks that required bots to handle more than single-turn conversations.
As the first attempt to support collaborative design of AI agents in community contexts, the current version of Botender focuses on enabling the design of single-turn, LLM-based conversational AI agents. This means that the bot and its underlying agents can process only one user message at a time and respond with a single message. Channel names are the only context information that allows the agents to tailor their behavior more specifically. Although participants were aware of this limitation from the beginning of the study, many expressed in the post-study survey a desire for expanded bot capabilities that would allow for greater customization to better address the unique needs of their communities. For example, participants shared feedback such as, “I would love for Botender to be able to take into context the full conversations!” (S15, G4), and “Being able to store data would be neat! For example, setting up a birthday reminder would require saving data about the different users” (S19, G1). They also expressed interest in additional bot actions beyond single-turn replies, such as “the ability to block messages” (S9, G6).
7.3.2. Opportunities: Future bots could build upon Botender’s system architecture to expand their capabilities for greater customization.
@@
As mentioned in Section 4.1, Botender’s system and agent architecture is designed to handle a broader range of platform events, actions, and context information. On the event side, in addition to listening to user messages, Botender’s listener can detect a wider range of platform events provided by the platform API, such as users joining channels, the creation of new threads, or changes to user permissions. These events can be translated into information that agents can interpret and use to provide appropriate action instructions. Similarly, on the action side, Botender’s action executor can carry out a broader range of actions available through the API, such as creating new channels, muting users when appropriate, or searching the internet for up-to-date information. [^16] We experimented with internet search, but found that it made case generation too slow and affected the user experience. It may become faster as LLMs improve. The context information available to agents can also be expanded depending on the event. For example, for the event of a user sending a message, the context could include the time, the user’s permission, a greater history of previous messages, or stored data about the community (Wang et al., 2025). Future systems could provide agents with richer context and expand their capabilities based on what would be most useful to communities. However, as prior research points out (Hwang et al., 2024), granting bots more permissions can inevitably raise concerns among community members about potential misuse and the risk of unrecoverable consequences. Striking the right balance between agent capabilities and user concerns will be an important challenge to address.